A convex lens those two faces has the same radius of curvature has a focal length of 16cm, refractive index of the glass is 1.2. find the radius curvature
1
Expert's answer
2019-12-19T11:39:55-0500

Optics


We need to find the Radius of the curvature.


Solution:


Given, focal length = "f = 16 \space cm"


Refractive index of the glass "= n = 1.2"


Let R be the radius of curvature of each face.


then, "R_1 = R\\

R_2 = - R"


We know the Lens marker's formula



"\frac {1}{f} =(n-1) [ \frac {1}{R_1} - \frac {1}{R_2}]"



Plug the values in this formula


"\frac {1}{16} =(1.2-1) [ \frac {1}{R} - \frac {1}{-R}]"



"\frac {1}{16} =(0.2) [ \frac {1}{R} + \frac {1}{R}]"




"\frac {1}{16} =(0.2) [ \frac {2}{R} ]"

"R = 0.2 \times 2 \times 16 \space cm"

"R = 6.4 \space cm"

Answer: Yes

Radius of the curvature = R = 6.4 cm